Find the derivatives of
1.y=3x^2-5x+5
2.f (x)= x+x^2÷2+x^3÷3
1
Expert's answer
2019-11-27T13:30:06-0500
"1) y' =(3x^2-5x+5)'= 3\cdot2\cdot x - 5 = 6x - 5;\\

2) f'(x)= (x+\frac{x^2}{2}+\frac{x^3}{3})'=1+\frac{2x}{2} + \frac{3x^2}{3}\\ = 1 + x + x^2"